# BAP v2: An Enhanced Task Framework for Instruction Following in Minecraft Dialogues

### v2

@@ 8.1 An additional key challenge for BAP evaluation
As shown in Figure 23, A initially directs B to place a red block on the left, but then corrects this to the analogous position on the right, requiring removal of the original block and placement of a new one. The baseline model correctly removes the original red block but fails to place a new one, instead removing the top green block randomly. The best model performs only the correct removal and places a new red block to the right, though not in the exact position specified—nonetheless, it is clearly a closer interpretation of A’s intent.
In this case, the F1 metric assigns both models a non-zero score of 0.5, as both have 1 true positive, 1 false positive, and 1 false negative. However, the type and color F1 metrics penalize the baseline model, scoring it 0.5 in each, while the best model achieves a perfect score of 1.
8.2 Evaluation on synthetic data
Going forward, evaluation on the synthetic data will also be important for making more meaningful progress on the BAP task. First, the more structured and simpler nature of the synthetic data allows us to assess basic model competencies. Second, as noted in Section 6.1.2, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. Evaluating on the synthetic data will help build models that do not show such counterintuitive behavior (Section 8.3 provides further evidence supporting this argument), and complements evaluation on $D_{\mathrm{mc}}$. In fact, future models should aim to surpass previous performance on $D_{\mathrm{mc}}$ while also achieving strong results on synthetic data, ideally improving on both.
8.3 Concurrent work
Concurrent to our work are the recently published (at EMNLP, Nov 12-16, 2024) works of Chaturvedi, Thompson, and Asher (2024) and Kranti, Hakimov, and Schlangen (2024), both exploring SOTA LLMs for the BAP task. The former finetunes the Llama-3-8B model (Dubey et al., 2024) on the original BAP data, resulting in the Nebula model, which predicts the next action sequence in a text-to-text fashion using all preceding conversation and action sequences as context. The latter employs in-context learning with GPT-4. Nebula achieves an F1 score of 39.2% and GPT-4 scores 39.0% (under the now legacy BAP evaluation). Nebula is marginally better, although both are very close and mark improved performance on the task.
As mentioned in Section 1, modeling is not the primary focus of our work. Instead, we focus on the BAP task as a whole, addressing key challenges in data and evaluation to propose BAP v2, an upgraded version of the task itself. We also illustrate the impact of this data on training a simple LLM and transformer-based model. This foundation enables future work to make more efficient and meaningful progress on the task, and sets the stage for more sophisticated models/LLMs. These works are therefore complementary to ours. A more detailed comparison to Chaturvedi, Thompson, and Asher (2024) is provided in the appendix (Section 15), given some other relevant aspects of their work.
Here, we evaluate Nebula using the v2 benchmark to further demonstrate the impact and robustness of our evaluation while providing a fairer assessment of its performance. We observe a 27.8% gain in performance (compared to the legacy evaluation), similar to the gain for our baseline $M_{\mathrm{mc}}$. We also discuss how the v2 evaluation now enables more accurate benchmarking of SOTA LLMs on the BAP task and highlights the significant room for improvement, further motivating the task itself. Additionally, it helps us confirm that spatial reasoning remains the key bottleneck for high performance, motivating future efforts leveraging our data and methods, which have shown effectiveness in addressing this challenge. We also evaluate Nebula on our synthetic data and find that, like the baseline $M_{\mathrm{mc}}$ (and the BERT-based $M_{\mathrm{mc}}^{\prime}$), it underperforms significantly despite being trained on the far more complex $D_{\mathrm{mc}}$. This suggests that models (including SOTA LLMs) trained solely on $D_{\mathrm{mc}}$ may not be as robust, reinforcing our argument of Section 8.2 about the importance of the synthetic data evaluation. Future work should explore finetuning models such as Llama-3-8B (Nebula) and newer LLMs using our synthetic data and training methodologies like CL.
BAP v2 evaluation
Similar to the 29.4% improvement we observed for our baseline model $M_{\mathrm{mc}}$ under the v2 evaluation on $D_{\mathrm{mc}}$, Nebula’s performance also increases from 39.2% to 50.1%, reflecting a 27.8% gain. This further underscores the impact and robustness of our v2 evaluation, offering a fairer reflection of Nebula’s performance. With this fairer evaluation now possible, we can more accurately gauge current performance on the BAP task and the remaining gap for improvement. While the 50.1% score marks improved performance, it also underscores that even with SOTA LLMs, significant room for improvement remains, thus reinforcing the task’s challenging nature and its importance for future research directions. Detailed tables for the full v2 benchmark are provided in the appendix (Tables 13 and 14). The observed trends generally align with those seen in our models such as the model’s relative strengths and weaknesses. They also shed some light on key areas for future improvement. For example, spatial reasoning (Location F1) is the key bottleneck for high performance on this task as noted in Section 4.4, and this continues to remain the case. Therefore, as demonstrated with $M_{\mathrm{mc}}$, our synthetic data and training methodologies can potentially help address this challenge, motivating future work.
Evaluation on the synthetic data
Table 9 shows Nebula’s performance (micro F1) on both the synthetic data ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) and the original $D_{\mathrm{mc}}$ test sets. Evaluating Nebula on the synthetic data reveals behavior consistent with our baseline model $M_{\mathrm{mc}}$ (and the BERT-based $M_{\mathrm{mc}}^{\prime}$), as reported in Section 6.1.2 and further discussed in Section 8.2 – it underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. This shows that models (including SOTA LLMs) trained solely on $D_{\mathrm{mc}}$ may not be as robust, and reinforces our argument from Section 8.2 about the importance of the synthetic data evaluation: it will help build models that don’t show such counter-intuitive behavior.
Model Test Data
$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
Nebula 51.5 33.1 30.9 50.1
Table 9: Nebula’s performance across datasets (and using the v2 evaluation on $D_{\mathrm{mc}}$)
8.4 Broader implications of our work
@@
While our focus in this work is on the Builder Action Prediction (BAP) subtask within the Minecraft Collaborative Building Task (MCBT) setting, the contributions have broader applications and takeaways as well. Below, we discuss their relevance for other subtasks within MCBT, for research building directly on the MCBT, and for similar domains and tasks in the field. This highlights the potential impact of our work and directions for future research.
@@
Although our primary focus is on BAP, the synthetic data and training methodologies introduced here are applicable to other MCBT subtasks constrained by the limited size of the MDC, such as the Architect Utterance Generation (AUG) subtask (Narayan-Chen, Jayannavar, and Hockenmaier, 2019) summarized briefly in Section 2.1.2. AUG and BAP represent two key pillars for building fully interactive Architect and Builder agents. Additionally, other subtasks, such as deciding when the Builder should act versus speak or generating clarification questions during dialogue, could also benefit from our synthetic datasets. Our simulators and data were carefully designed to emulate the nuances of the MCBT itself (Section 5.1), making them directly usable or adaptable to these subtasks. The simulators’ parameterized nature enables generating tailored data distributions to this end, supporting further advancements across the broader MCBT ecosystem.
@@
Potential applications of our work also go beyond the MCBT. Several works directly build upon or are inspired by the MCBT/MDC, including IGLU (Mohanty et al., 2024), Bonn et al. (2020), Thompson, Hunter, and Asher (2024), and Bonial et al. (2021). Therefore, future work on them can also potentially benefit from the synthetic data.
@@
Our findings also offer broader takeaways for research on tasks involving grounded, multimodal task-oriented dialogue systems operating in dynamic, spatially complex environments. First, in many realistic scenarios, data collection for complex situated dialogue tasks is often impractical or expensive. Our work illustrates that synthetic data generation can be a viable and scalable alternative to simply collecting more data. Second, we demonstrate that rich synthetic data can be generated for such tasks in game-based or simulated environments like Minecraft by carefully emulating task and game-specific nuances. This approach produces full embodied task oriented dialogues combining utterances, environment actions, and other game state information, offering a viable strategy for similar game-based/simulated and grounded dialog tasks –– an underexplored area as highlighted in Section 9.2. Finally, techniques like curriculum learning demonstrate potential and can be explored further in similar domains/tasks in combination with synthetic data.
9 Related Work
9.1 Task and Dataset
We discuss related work to illustrate that the MDC and BAP remain a challenging and relevant testbed for grounded language instruction following.
Vision-and-Language Navigation (VLN) (Anderson et al., 2018), and its dialog counterpart, Cooperative Vision-and-Dialog Navigation (CVDN) (Thomason et al., 2019), explore instruction following in photorealistic navigation settings. However, Minecraft dialogues are more complex due to their asynchronous communication, longer turn sequences, and dynamic environment changes during construction. Unlike navigation, where referring expressions target existing objects, construction often involves instructions for objects yet to be built. And although more recent navigation tasks require real vision, their underlying world state space (as defined by fixed viewpoints and the underlying navigation graph) is just as highly discretized. Our task does not require vision, but poses an arguably more challenging planning problem, since its action space is much larger (7623 possible actions vs. six actions in the vision-language navigation work). Tasks involving vision and language-based task completion beyond navigation have also been explored. ALFRED (Shridhar et al., 2020) is a benchmark for mapping natural language instructions and egocentric vision to action sequences for household tasks in simulated environments. However, similar to VLN and unlike the MDC, it lacks dialogue and relies solely on planner-based demonstrations instead of human ones. Efforts to combine vision and dialogue navigation with task completion include CEREALBAR (Suhr et al., 2019), which operates in a game-based setting. Yet, similar to CVDN, it uses turn-based dialogue, contrasting with the free-form, asynchronous communication of the MDC.
This limitation is addressed by the more recent benchmark TEACh (Padmakumar et al., 2022), which includes free-form dialogue, similar to the MDC, and features over 3000 human-human interactive dialogues for completing household tasks in simulation and with object state changes. However, unlike the MDC, the instruction giver in TEACh is provided with a predefined plan of steps, while the Architect in the MDC independently devises a unique plan for constructing the target structure. This results in greater creativity and diversity in the MDC, with multiple possible plans for the same structure. Although the TEACh dataset is larger (by number of #dialogs), its dialogues and utterances are shorter. Additionally, the MDC allows for greater abstraction in instructions, such as referring to substructures, individual blocks, or shapes, whereas TEACh focuses on well-defined objects in the environment. Similar to VLN and CVDN, referring expressions in TEACh refer to objects that exist in the world, but construction instructions like those in the MDC frequently refer to objects that need to be built.
Among Minecraft-related efforts, IGLU is most closely related to the MDC/MCBT and was directly inspired by it. The IGLU dataset, unlike the MDC, is single turn only, and not dialogue based. It comprises 8136 instruction-action pairs, including 1056 ambiguous instructions with corresponding clarification questions. The task setting is similar to the MCBT but incorporates some simplifications which make it relatively less challenging. E.g., cardinal directions (e.g., North, South) are displayed in the voxel world, which simplifies instruction giving/following through the use of absolute rather than relative references (e.g., left, right). IGLU also introduces a very small multi-turn dataset (127 dialogues) based on this same task setting. It features strict turn-taking and clearly demarcated instructions and clarification questions. This contrasts with the MDC’s free-form, asynchronous dialogues, which are less structured and whose utterances have more diversity.
Some other works also build on or are inspired by the MCBT/MDC. For instance, Bonn et al. (2020), Thompson, Hunter, and Asher (2024), and Bonial et al. (2021) provide additional linguistic annotations for the MDC.
9.2 Synthetic Training Data
Synthetic data has proven beneficial for various tasks, including those mentioned in Section 9.1. VLN agents, constrained by limited human instruction data and diversity in training environments, often struggle with complex language grounding and spatial language understanding. Kamath et al. (2023) address this by using synthetic data, achieving SOTA performance on the RxR dataset (Ku et al., 2020), while Wang et al. (2023) are similar in spirit and achieve SOTA on the aforementioned CVDN and other benchmarks. Similarly, Kang et al. (2023) demonstrate strong performance gains in low-data regimes for Visual Dialog Das et al. (2017) by leveraging synthetic data. Synthetic data has also been effective for dialogue systems, particularly those focused on text-only dialogues (Kim et al., 2022; Bao et al., 2023; Zhan et al., 2023). Zhan et al. (2024) is an example of extending this to multi-modal dialogue agents, demonstrating the effectiveness of synthetic visual descriptions in enhancing agents’ grounding capabilities.
However, fewer efforts have addressed synthesizing task-oriented embodied dialogues (involving both utterances and environment actions). Padmakumar et al. (2023) were the first to design a framework to do so. They extend agenda-based dialogue simulation (Schatzmann and Young, 2009) to a multimodal embodied agent framework, and demonstrate the impact of the synthetic dialogues on the TEACh task (see Section 9.1 for a comparison of TEACh to MDC). Our work serves to provide another useful example in this general domain, further advancing synthetic data generation for such complex tasks. Dan, Han, and Roth (2021) is another closely related work that demonstrates the potential of synthetic data in the Blocks World domain. They use simpler synthetic data for the benchmarks of Bisk, Yuret, and Marcu (2016), which require understanding single-shot instructions that transform one world state to another using simulated 3D blocks. Blocks are viewed from a fixed bird’s eye perspective, initialized randomly in the initial world state, and uniquely identifiable. The varying Builder perspective and lack of easily identifiable referents, along with the need to understand utterances in a dialogue context, make the BAP task a much more challenging problem.  [^7] Also, unlike traditional Blocks World, Minecraft allows blocks to float (requiring non-monotonic action sequences where placement is followed by removal), or attach to any side of an existing block. Our work on synthetic data thus serves as a demonstration of the potential for richer, embodied task-oriented dialogue tasks that extend beyond simpler Blocks World settings.
10 Conclusion and Future Work
In this work, we revisited the challenging Builder Action Prediction (BAP) subtask within the Minecraft Collaborative Building Task (MCBT) setting, addressing key shortcomings in evaluation and training data. We introduced BAP v2, an upgraded version of the task, comprising two core contributions: (1) an enhanced evaluation benchmark that includes a cleaner test set and fairer, more insightful metrics, and (2) diverse synthetic training datasets generated from novel Minecraft dialogue and target structure simulators. These contributions together provide a robust foundation for future research on the BAP task. Additionally, our synthetic data demonstrates the potential to train more performant and robust models, even with relatively simple training methods and architectures. Although modeling was not the primary focus of this work, we also illustrated the impact of the synthetic data and training methodologies on training a simple LLM and transformer-based model, thus validating the robustness of our approach, and setting the stage for more sophisticated models/LLMs going forward.
Future work can therefore build on our work to explore more advanced LLMs and architectures while leveraging the BAP v2 framework, alongside training methodologies such as Curriculum Learning (CL). Further improvements could stem from experimenting with more sophisticated CL approaches or entirely new training paradigms to get even more out of the synthetic data. Additionally, as highlighted in Section 8.1, BAP evaluation remains a nuanced challenge, warranting deeper investigation. Finally, Section 8.4 outlined the broader implications of our work, while also highlighting potential directions for future research.

### v3

@@ 7.2 Qualitative evaluation
    Example E
    Figure 12: Auxiliary metrics provide further insight than the overall F1 score (Scenario 2)
    Scenario 3: Mistakes that warrant a more semantic evaluation
    Example F
    Example G
    Figure 13: Examples that warrant a more semantic evaluation (Scenario 3)
    We now present three different scenarios, i.e. types of examples of output from Llama-CRAFTS and Nebula that, on the one hand, illustrate different ways in which Llama-CRAFTS arguably displays a better spatial understanding, and, on the other hand, allow us to discuss how this is reflected (or not) in our various metrics.
    Scenario 1: The overall F1 score is sufficient when one model is correct and the other makes mistakes
    Figure 11 shows three examples that Llama-CRAFTS gets completely right (receiving perfect scores across all metrics), and Nebula does not. In Example A, A instructs B to remove the four corners of the orange ring. Nebula removes many extraneous blocks but only three corners (achieving an overall F1 of 0.38). In Example B, A asks B to add another red block one tile below and to the left of the last block B had placed. Nebula only correctly identifies that a single red block should be placed below the last one, but goes too far down (2 tiles), and misses the second spatial relation (left) altogether, resulting in a perfect type and color F1 scores of 1.0. and an overall F1 of 0.0. In Example C, A asks B to “copy" the red ring in blue. Instructing Builders to copy substructures is a natural and efficient strategy that the human Architects in the MDC employ frequently. Nebula achieves perfect type and color scores, but misplaces one of five blocks, resulting in location and shape scores of 0.8. In these clear cases, the strict F1 metric is obviously sufficient to capture the qualitative difference between the outputs.
    Scenario 2: Both models make mistakes, but the auxiliary metrics provide further insight
    Figure 12 shows examples where both models make mistakes (although Llama-CRAFTS’s mistakes are arguably less egregious), and illustrates how our metrics capture different aspects of this task. In Example D, B is asked to make “a red 3 x 3 square" that is “vertical" (and presumably includes the four red blocks already placed). Both models add eight red blocks to create squares that leave out the bottom three red blocks and only include the red block above. Since Llama-CRAFTS’s square is actually vertical, it achieves a score of 0.77 on all metrics, whereas Nebula built a horizontal square, and hence achieves overall, shape and location scores of 0.31 (and type and color scores of 0.77). That is, although the overall score correctly identifies Llama-CRAFTS as the better model, it is the auxiliary metrics that provide additional insight and assign credit to Nebula. In Example E, B has just finished the blue-purple-orange (from right to left) line, and A references earlier instructions when saying that “the other one: [should be like the one B already made but with] green red yellow [instead of blue purple orange] …". Based on the earlier dialogue (also not shown in the figure), this second line should include the green block already on the board. [^10] This example illustrates a shortcoming of our figures: although BAP models have access to the entire preceding context, the reader does not. Here, both models receive a 0.0 overall F1 score, Nebula because it produces no net actions (receiving 0.0 scores on all metrics), and Llama-CRAFTS because none of its blocks appear in the intended places. However, because Llama-CRAFTS actually places a diagonal line of the correct color (albeit on the other side of the previous line and with the pattern starting from the left instead of the right), it receives type, color, and shape scores of 0.88. In this case, the auxiliary metrics are necessary to identify that Llama-CRAFTS’s output is actually quite similar to the desired structure.
    Scenario 3: Shortcomings of our metrics
    Figure 13 illustrates shortcomings of our automatic evaluation procedures, and suggests the need for a more complex, semantic evaluation. In Example F, B is asked to create a (diagonal) line of four blocks inside the ring-shaped structure, “from top left to bottom right". Nebula places three disconnected blocks inside the structure (achieving overall and location scores of 0.0, type and color scores of 0.86, and a shape score of 0.57), while Llama-CRAFTS places a (horizontal) line of four blocks across the inside top edge, as well as an additional block in the (outside) bottom right corner (achieving overall, location and shape scores of 0.22, and type and color scores of 0.89). The fact that Nebula’s shape score is higher than Llama-CRAFTS’s points to a limitation of our shape metric (two of Nebula’s blocks can be mapped to the intended diagonal) This example also suggests a need for a more semantic evaluation: none of our metrics capture the fact that, arguably, Llama-CRAFTS showed at least a partial understanding of the “line" concept (and, in a way, even the “bottom right"), but Nebula did not. In Example G, A asks B to place two blocks on the “"other side"" of “"what you just made"" (i.e. the line of four blocks on the board). Such implicit references to spatial layouts and to existing structures are very common in the MDC dialogues. But we see another shortcoming of our metrics: because both models place a row of two green blocks on the ground (and because our shape metric allows for rotations on the horizontal plane), they achieve location, shape and overall F1 scores of 0.5, and type and color scores of 1.0, even though Llama-CRAFTS arguably understood “other side" and “90 degrees" better than Nebula.
    7.3 Takeaways
    Our qualitative analysis is consistent with the quantitative results in Table 7: Both models are adept at identifying simpler aspects of the task, such as action type and color, reflected in their high Type and Color F1 scores. Llama-CRAFTS also demonstrates a superior understanding of required block shapes (e.g., diagonals, corners), which aligns with its higher Shape F1 score. The primary differentiator, however, is spatial reasoning. Llama-CRAFTS’s improvement in this area, demonstrated by its higher Location F1 score, validates the impact of our synthetic data and its richer input representations.
    Error Analysis
    However, our error analysis illustrates that Llama-CRAFTS still makes significant errors in precise spatial understanding, indicating that much work remains to be done to solve this task.
    Evaluation Challenges and the Role of Auxiliary Metrics
    The qualitative analysis reveals a more nuanced challenge for fair evaluation: the overall F1 metric lacks the sensitivity to distinguish differences in error quality. As our qualitative analysis reveals, it is possible for two models to make different types of errors, with one being qualitatively "better" or closer to the correct solution, yet the F1 metric may not capture this distinction. This demonstrates that our proposed auxiliary metrics can play a crucial role, either by identifying the types of mistakes a model makes or by identifying cases where the model displayed at least a partial understanding of the instruction. At the same time, our analysis underscores the inherent challenges in automated evaluation for the BAP task.
    The examples in Scenario 3 also illustrate the difficulty of creating a clear, unambiguous rubric for grading model output on an absolute scale, making human evaluation equally complex. This complicates not only absolute scoring but also comparative pairwise evaluations (i.e., "Is model A better than model B?"). While our qualitative analysis includes examples where a clear distinction can be made, many cases are far more ambiguous. In fact, we initially considered human evaluation but ultimately did not pursue it due to these challenges. Also, given that the primary goal of this work is to provide a holistic revision of the entire BAP framework—spanning evaluation, data, and modeling—we concluded that a deep investigation into human evaluation methodologies was beyond the current scope and needed thorough investigation owing to the aforementioned challenges. This highlights the need for further research into BAP evaluation by building on our v2 benchmark.
    Implications for LLMs
    Finally, the persistence of these subtle spatial errors, as observed above in Section 7.2, suggests limitations in how current text-only LLMs handle fine-grained 3D spatial grounding in embodied tasks, specifically in a 3D, embodied, collaborative dialogue setting (see Section 7.5 below for more discussion in the broader context of related work), thus motivating future work.
    7.4 Related Work: BAP evaluation
    Chaturvedi, Thompson, and Asher (2024) also identify some of the issues with BAP evaluation on $D_{\mathrm{mc}}$ that we do, and hence propose evaluation on synthetic data. They design simpler, non-dialogue-based data with corresponding evaluation metrics. The data is single-turn and evaluates a model’s ability to follow an instruction in two scenarios: when the board is empty and the instruction is to build a single shape, or when a single shape is already present and the instruction involves placing or removing one block.
    While this data is intentionally simple to test elementary Builder capabilities, it overly simplifies the task and therefore deviates significantly from the target MCBT/BAP. It lacks a dialogue component, game history, and relative spatial references such as "left" and "right" as the Builder’s position and orientation are not considered. The world state/built structure is also fairly minimal, either empty or containing a single shape. While useful for elementary testing, it is a bit unclear how well insights from this data translate to the full BAP task. In contrast, our synthetic data remains simpler than $D_{\mathrm{mc}}$ while also closely emulating the MCBT, addressing these limitations and preserving many of the key BAP features and challenges, such as diverse instructions, rich spatial relations and floating blocks for example (more in Sections 2.1.1 and 5). Although designed/motivated for training more than evaluation, its simplicity allows assessment of basic model competencies, while its emulation of the MCBT enhances its applicability to the overall BAP task, thus striking a better balance from an evaluation standpoint. Section 6.3 details additional benefits of evaluating on our synthetic data.
    Additionally, our synthetic data evaluation, combined with fairer F1, enables meaningful performance comparisons between synthetic data and $D_{\mathrm{mc}}$. Chaturvedi, Thompson, and Asher (2024) are unable to conduct such comparisons (albeit understandably) due to the limitations of the legacy/original F1 metric.
    7.5 Related Work: Spatial Reasoning
    Early efforts to understand spatial reasoning in language models focused on extracting spatial relations from text and observing what conclusions about a real world a model trained only on rich language data could learn (Mirzaee and Kordjamshidi, 2022). Spatial reasoning was often explored in simple text-based games (Shi, Zhang, and Lipani, 2022) or with carefully constructed reasoning tasks (Rizvi, Zhu, and Gurevych, 2024).
    Spatial reasoning has also been explored in conjunction with other modalities. In Vision-and-Language Navigation many explore it over scene-graphs (Gopinathan et al., 2024; Li et al., 2024a). Others explore spatial reasoning with text and image inputs. Bisk et al. (2018) explore what vision and language activations models learn from block tasks represented in 2D images: mapping out how models interpreted spatial relations but also verbs like "mirror" or "twist". Du et al. (2024) also explore how VLMs can reason over space in photographs, but in reference to objects within those objects (i.e. an object being on the right of the photo but to the left of a person in the photo). However, Kamath, Hessel, and Chang (2023) found that VLMs spatial reasoning can lag well behind human performance on VQA-style tasks and Wang et al. (2024) find that VLMs often do not make adequate use of the images that they are given, especially when provided with rich text representations.
    Unfortunately, there is not much parallel work exploring the spatial capabilities of LLMs on text-only inputs, particularly for embodied tasks. Yamada et al. (2024) explore some basic grid navigation abilities of prompted LLMs, which struggle to traverse simple grids represented in text. Aghzal, Plaku, and Yao (2024) consider the problem from the perspective of text-only grid path-planning, showing that models can effectively link coordinates and directions, but struggle to generalize to larger grids with more obstacles. Others have explored LLMs in playing chess (Feng et al., 2023a; Zhang et al., 2025b) or sliding puzzle games (Chollet et al., 2025) to some degree of success, but we found no works combining 3D text-only grids and embodied collaborative dialogue. Together, these gaps in the literature motivate the MDC setup: by excluding the visual modality used in other tasks, it allows us to explore the baseline spatial capabilities of text-only LLMs on embodied tasks in a small controlled setting that is still generalizable to concrete end-goals.
    8 Future Work
    We now discuss the implications of our work and outline directions for future research. This discussion is organized into two parts: the path forward for BAP (Section 8.1), and the broader impact of our work beyond BAP (Section 8.2).
    8.1 The path forward for BAP
    The introduction of the BAP v2 framework, while a significant step forward, also illuminates several key directions for future research. The work detailed in this paper suggests that the next advancements on the BAP task can be organized around three core pillars: evaluation, data, and modeling.
    Advancing Evaluation
    Section 7.3 discusses the persisting problems around both automated and human evaluation for BAP, in spite of the progress made in BAP v2. This highlights the need for further research into evaluation by building on our v2 benchmark.
    Enhancing Synthetic Data Generation
    We note the broader challenges in using LLMs to generate coherent, state-aware synthetic data for embodied dialogue tasks (Section 5.6). A promising research direction is to explore methods for overcoming these limitations, potentially leveraging LLMs to create even more diverse, complex, and naturalistic training dialogues that still maintain grounding in the 3D world state.
    Improving Model Capabilities
    With a new SOTA F1 score of 53.0, there is still considerable room for improvement. Key areas for future modeling work include
    • Spatial Reasoning: Our analysis (Section 7) confirmed that highly accurate and fine-grained 3D spatial reasoning remains a primary bottleneck, even for top-performing models. Future models must develop more robust mechanisms for interpreting complex, perspective-dependent spatial language.
    • Advanced Training Regimes: More sophisticated training strategies could unlock further gains. For instance, exploring curriculum learning, where models are first trained on simpler synthetic data before moving to the more complex human-generated dialogues, could prove effective.
    • Architectural Exploration: The relative success of a smaller fine-tuned specialist model motivates deeper exploration of this paradigm. Future efforts can build on Llama-CRAFTS with more advanced architectures or fine-tuning techniques. Concurrently, the synthetic data could also be leveraged to improve the ICL performance of larger, general-purpose models (Section 6.4).
    8.2 Broader implications of our work beyond BAP
@@
    While our focus in this work is on the BAP subtask within the MCBT setting, our contributions have broader applications and takeaways as well. Below, we discuss their relevance for other subtasks within MCBT, for research building directly on the MCBT, and for similar domains and tasks in the field; thus highlighting the potential applications of our work and directions for future research.
@@
    Although our primary focus is on BAP, the synthetic data and training methodologies introduced here are applicable to other MCBT subtasks constrained by the limited size of the MDC, such as the Architect Utterance Generation (AUG) subtask (Narayan-Chen, Jayannavar, and Hockenmaier, 2019) summarized briefly in Section 2.1.2. AUG and BAP represent two key pillars for building fully interactive Architect and Builder agents. Additionally, other subtasks, such as deciding when the Builder should act versus speak or generating clarification questions during dialogue, could also benefit from our synthetic datasets. Our simulators and data were carefully designed to emulate the nuances of the MCBT itself (Section 5), making them directly usable or adaptable to these subtasks. The simulators’ parameterized nature enables generating tailored data distributions to this end, supporting further advancements across the broader MCBT ecosystem.
@@
    Potential applications of our work also go beyond the MCBT. Several works directly build upon or are inspired by the MCBT/MDC, including IGLU (Mohanty et al., 2024), Bonn et al. (2020), Thompson, Hunter, and Asher (2024), Bonial et al. (2021), and Madge et al. (2025). Therefore, future work on them can also potentially benefit from the synthetic data.
@@
    Our findings also offer a key takeaway for research on embodied, multimodal, task-oriented dialogue in dynamic and spatially complex settings. We demonstrate a methodology for generating rich synthetic data within game-based or simulated environments by carefully emulating task- and game-specific nuances. This approach yields complete embodied task-oriented dialogues that integrate utterances, environmental actions, and other game state information, providing a robust strategy for such complex dialogue settings—an area that remains underexplored, as highlighted in Section 5.6. This enables future work in the domain.
    9 Conclusion
    In this work, we addressed the challenges of the Builder Action Prediction (BAP) task, a demanding testbed for grounded instruction following within the Minecraft Collaborative Building Task. To foster more efficient and meaningful progress, we introduced BAP v2, a holistically re-examined and upgraded task framework. BAP v2 is built on three core contributions. First, we established an enhanced evaluation benchmark, featuring a cleaner test set and fairer, more insightful metrics. This new benchmark not only provides a more accurate measure of performance but also revealed that spatial reasoning is the primary performance bottleneck. Second, to address both data scarcity and the spatial reasoning challenge, we developed novel dialogue simulators to generate diverse synthetic training data rich in spatial language. Third, we leveraged this new data to train stronger models. Our new SOTA model, Llama-CRAFTS, utilizes rich, context-aware input representations and the synthetic data to achieve an F1 score of 53.0 —a significant 6 point improvement over previous work.
    While this result marks a substantial advance, it also underscores the task’s remaining difficulty, even for SOTA LLMs. By systematically improving the evaluation, data, and modeling paradigms, the BAP v2 framework provides a robust and reliable foundation for future research. It paves the way for developing more sophisticated models with advanced training regimes, creating richer synthetic data, and exploring new evaluation techniques to push the boundaries of what is possible for interactive agents in complex, collaborative, and embodied environments.